OpenStudy (purplemexican):

Using the quadratic formula to solve 2x2 = 4x – 7, what are the values of x?

OpenStudy (lynfran):

rewrite the equation to get 2x^2-4x+7=0 where a=2 , b=-4 and c=7 now substitute the values in the quadratic formula \[\frac{ -b \pm \sqrt{(b ^{2}-4ac)} }{ 2a }\]

OpenStudy (lynfran):

\[\frac{ -(-4)\pm \sqrt{((-4)^{2}-4(2)(7))} }{ 2(2) }\]now simplify...

OpenStudy (purplemexican):

its been a while since ive done this and ive lost it

OpenStudy (anonymous):

Linda I need you for a minute

OpenStudy (anonymous):

continuing the step: \[\frac{ (4)\pm \sqrt{(16)-(56)} }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ (4)\pm \sqrt{(-40)} }{ (4) }=\frac{ (4)\pm \sqrt{(40)(-1)} }{ 4 }\]

OpenStudy (anonymous):

complex numbers: \[\frac{ (4)\pm(\sqrt{40})(\sqrt{-1}) }{ 4 }=\frac{ (4)\pm(\sqrt{40})i }{ 4 }\] :)